Ibuprofen, a headache remedy, contains 75.69% C, 8.80% H, and 15.51% O by mass, and has a molar mass of 206 g/mol.

Express your answers as chemical formulas separated by a comma.
Answer #1

moles C = 75.69 g /12.011 g/mol = 6.302
Moles H = 8.80 g / 1.008 g/mol =8.73
Moles = = 15.51 g / 15.9994 g/mol = 0.9694

we divide by the smallest number
6.302/ 0.9694 = 6.5 =>C
8.73 / 0.9694 = 9 => H
0.9694 / 0.9694 = 1 => O

to grt whole numbers we multiply by 2

C13 H18O2 is the empirical formula ( MM = 206 g/mol)